OpenStudy (anonymous):

Solve the equation. |x – 3| = 8 A. {–11, –5} B. {–11, 5} C. {5, 11} D. {–5, 11} @ganeshie8 @bigblock557 @Zale101 @raffle_snaffle

OpenStudy (anonymous):

are you sure?

OpenStudy (ash2326):

@adamarbour Please don't just handover the answer, please explain the working.

OpenStudy (anonymous):

oh this is how your solve it x-3=8 x-3=-8 x=11 x=-5

OpenStudy (ash2326):

@TakeA.I.M. Do you follow this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @adamarbour oh this is how your solve it x-3=8 x-3=-8 x=11 x=-5 \(\color{blue}{\text{End of Quote}}\) This is true, but I like to start with: x-3=8 -(x-3)=8 This is consistent with inequalities.